OpenStudy (anonymous):

If I have a quadratic law: y = (ax^2)-(bx)+c; and I do have several values of x and y, say 10 values, then how do I find the coefficients, a,b,c ? I want to do it using MATLAB.

OpenStudy (anonymous):

You get a system of linear equations with a,b,c ... Suppose (1,2),(4,5),(6,7) are 3 points on the curve. from y = (ax^2)-(bx)+c plug in these values of (x,y) to get 1=4a-2b+c 4=25a-5b+c 6=49a-7b+c Now you solve the system to get (a,b,c). Hence you'll usually need three sets of values to determine (a,b,c)